OpenStudy (anonymous):

Nathan works part-time from home as a data entry professional. He enters data in forms and is paid based on the entry accuracy rate. The table below shows his compensation. Based on his regular performance, the equation below can be used to determine C, the total amount Nathan earns a month entering data, where x represents the total number of forms he completes that month. C = 18(0.7x) + 12(0.2x) + 0(0.1x) What is the minimum number of forms he must complete in a month to earn a total of at least $1,650?

OpenStudy (anonymous):

so C = 1650 1650 = 18(0.7x) + 12(0.2x) + 0(0.1x) solve for x: 1650 = 12.6x + 2.4x + 0 1650 = 15x x = 110 forms